Find The equation of the parabola whose focus is (5,3)and directrix is x+5=0

Grade:12

1 Answers

Arun
6 years ago
as we know parabola is a locus of point in which distance of it from the focus is same the perpendicular distance from directrix
hence let that point is (x,y)
now
\sqrt(x-5)2 + (y-3)2 = x+5
x^2+ 25 – 10x + y^2 +9 – 6y = x^2 + 25+ 10x
y^2 – 6y = 20x + 16